The number of visits to a website in one hour has the following probability mass function

{1/10 , 𝑥=0

𝑝(𝑥)={ 𝑘𝑥/10 , 𝑥=1,2,3

{𝑘(6−𝑥)15, 𝑥=4,5


a) Determine the value of 𝑘.


With the value of 𝑘 obtained:

b) Find the probability of having

i) 3 visits in one hour.

ii) less than 2 visits in one hour.

iii) more than 6 visits in two hours.


c) By using central limit theorem, estimate the probability of having less than 253 visits in 100 hours.


1
Expert's answer
2020-08-04T19:10:30-0400

Solution a)


"P(x0) + P(x1) + P(x2) + P(x3) + P(x4) + P(x5) = 1"


"P(x) = \frac{1}{10}" for x = 0

"P(0) = \frac{1}{10}"


"P(x) = \frac{kx}{10}" for x = 1, 2, 3

"P(1<= x <= 3) = \frac{k}{10} + \frac{2k}{10} + \frac{3k}{10} = \frac{6k}{10}"


"P(x) = k(6-x)" for x = 4, 5

"P(4<= x <= 5) = k(6-4) + k(6-5) = 3k"


"1 = \frac{1}{10} + \frac{6k}{10} + 3k"

"1 - \frac{1}{10} = \frac{9}{10} = \frac{6k + 30k}{10} = \frac{36k}{10}"

"(9 * 10) = (36k * 10)"


Answer: 

"K = \frac{90}{360} = 0.25"


Solution b)

Part i) 3 visits in 1 hour


"P(x=3) = \frac{0.25x}{10} = \frac{0.25* 3}{10} = 0.075"

Answer: 0.075


Part ii) less than 2 visits in 1 hour

"P((x = 0 \, or \,x=1) = \frac{1}{10} + \frac{0.25}{10} = 0.125"

Answer: 0.125


Part iii) more than 6 visits in 2 hours

Let x be the visits in the first hour and y be visits in the second hour


"Max(x+y) = 10"

"Min(x+y) = 7"


"P(x + y > 6) = P(x + y = 7 \, or \,x + y = 8 \\\, or \,x + y = 9\, or \, x + y = 10 \, or \,x + y = 11\, or \,x + y = 12)"


"P(x + y = 7) = P(x = 5 \,and \, y = 2) + P(x = 4 \, and \, y = 3) + P(x = 3 \,and \, y = 4) + P(x = 2 \,and \,y = 5)"

"= (0.25(6-5) * \frac{0.25 * 2}{10} + 0.25(6-4) * \frac{0.25 * 3}{10} + \frac{0.25 * 3}{10} * 0.25(6-4) + \frac{0.25 * 2}{10} * (0.25(6-5)"

= 0.1


"P(x + y = 8) = P(x = 5 \, and \,y = 3) + P(x = 4 \,and \,y = 4) + P(x = 3 \,and \,y = 5)"

"= (0.25(6-5) * \frac{0.25 * 3}{10} + 0.25(6-4) * 0.25(6-4) + \frac{0.25 * 3}{10} * (0.25(6-5)"

= 0.2875


"P(x + y = 9) = P(x = 5 \, and \,y = 4)+ P(x = 4 \, and \,y = 5)"

"= (0.25(6-5) * (0.25(6-4) + (0.25(6-4) * (0.25(6-5)"

= 0.125


"P(x + y = 10) = P(x = 5 \, and \, y = 5)"

"= (0.25(6-5) * (0.25(6-5)"

= 0.0625


"P(x + y = 11) = 0"

"P(x + y = 12) = 0"


"P(x + y > 6) = 0.1 + 0.2875 + 0.125 + 0.0625"

Answer: 0.575


Part c) Probability of having less than 253 visits in 100 hours

Average visits per hour (x) = "\frac{253}{100}" = 2.53

Expectation of x = "\mu = \sum{x p(x)}"


X            px                      xpx                        p(x) * (x-\mu)^2

0             "\frac{1}{10}"                       0                        1.296

1            "\frac{0.25}{10}"                   0.025                     0.169

2            "\frac{0.5}{10}"                     0.1                         0.128

3             "\frac{0.75}{10}"                  0.225                     0.027

4             0.5                        2                         0.08

5             0.25                   1.25                       0.49


"\mu = 0 + 0.025 + 0.1 + 0.225 + 2 + 1.25 = 3.6"


"Var(x) = \sum{ p(x) * (x-\mu)^2} = 2.433"

"\sigma = \sqrt{2.433} = 1.5598"


"Z value = \frac{x- \mu}{sd} = \frac{2.53- 3.6}{1.5598} = -0.686"


At z value "= -0.686" ,"p = 0.2464"


Answer: 0.2464